OpenStudy (anonymous):

Simplify: i4 + i3 + i2 + i

terenzreignz (terenzreignz):

\[\Large i=\sqrt{-1}\]assuming?

OpenStudy (anonymous):

are these exponents?

OpenStudy (anonymous):

\[i^4 + i^3 + i^2 + i\]

OpenStudy (anonymous):

satellite, what is the trend for i? i forgot and dont have my notes, i know it goes in set of 4, so 1i, 2i, 3i,4i then repeats. so would you know what im speaking about?

OpenStudy (anonymous):

no, no exponents, The answer choices are A) 0 B) 1 C) -1 D) i

OpenStudy (anonymous):

then i am going to guess they are exponents in fact i am sure they are

OpenStudy (anonymous):

@sjerman1 the pattern is \[i^0=1, i^1=i, i^2=-1, i^3=-i, i^4=1, ...\]

OpenStudy (anonymous):

*YES! sorry about that. They're exponents.

OpenStudy (anonymous):

thank you, this will help @Charles10 quite a bit. @Charles10 WRITE THAT DOWN

OpenStudy (anonymous):

@Charles10 compute as i wrote above

OpenStudy (anonymous):

i did, i appreciate it. @satellite73 @sjerman1

OpenStudy (anonymous):

They are exponents, you are @satellite73 Remember, complex numbers with imaginary parts are added the same way complex numbers without imaginary parts (Real Numbers). Similarly, imaginary parts are raised to powers the same way. This gives:\[i = \sqrt{-1} \rightarrow i^4 +i^3+i^2+i=(\sqrt{-1})^4+(\sqrt{-1})^3+(\sqrt{-1})^2+\sqrt{-1}\]\[(-1)(-1)+(-1\sqrt{-1})+(-1)+\sqrt{-1}=1-\sqrt{-1}-1+\sqrt{-1}=0\]Get it? @Charles10

OpenStudy (anonymous):

They are exponents, you are @satellite73 Remember, complex numbers with imaginary parts are added the same way complex numbers without imaginary parts (Real Numbers). Similarly, imaginary parts are raised to powers the same way. This gives:\[i = \sqrt{-1} \rightarrow i^4 +i^3+i^2+i=(\sqrt{-1})^4+(\sqrt{-1})^3+(\sqrt{-1})^2+\sqrt{-1}\]\[(-1)(-1)+(-1\sqrt{-1})+(-1)+\sqrt{-1}=1-\sqrt{-1}-1+\sqrt{-1}=0\]Get it? @Charles10

OpenStudy (anonymous):

They are exponents, you are @satellite73 Remember, complex numbers with imaginary parts are added the same way complex numbers without imaginary parts (Real Numbers). Similarly, imaginary parts are raised to powers the same way. This gives:\[i = \sqrt{-1} \rightarrow i^4 +i^3+i^2+i=(\sqrt{-1})^4+(\sqrt{-1})^3+(\sqrt{-1})^2+\sqrt{-1}\]\[(-1)(-1)+(-1\sqrt{-1})+(-1)+\sqrt{-1}=1-\sqrt{-1}-1+\sqrt{-1}=0\]Get it? @Charles10